# Convergence analysis of Levenberg-Marquardt method with Singular Scaling for non-zero residue nonlinear least-squares problems [^1] This work is supported by FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360]

### v1

@@ 4.2 Diminishing Rank
Since the matrix rank is a lower semicontinuous function, for a given sequence $\{x_{k}\}$ converging to $x^{*}\in X^{*}$ it may happen that $\text{rank}J(x_{k})>\text{rank}J(x^{*})$, for every $k$ sufficiently large. We refer to this situation as “diminishing rank” scenario. In this case the convergence results will depend on the specific choice $\lambda_{k}=\|J(x_{k})^{T}F(x_{k})\|^{r^{\prime}}$, with $r^{\prime}\leq r$ from Assumption 4. To simplify the exposition, we consider $r^{\prime}=r$ in the results that follow.
Lemma 10.
Suppose that Assumptions 1-3 and 4 are valid, and $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$. If $x_{k},x_{k+1}\in B(x^{*},\delta/2)$ and $\textup{dist}(x_{k},X^{*})<\varepsilon$, where
$\varepsilon\leq\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r}$ (47)
with $\eta\in(0,1)$, then
@@
Proof.
Note that from Assumption 4 and Lemma 6, for $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, we have $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$, with
$c_{1}=\frac{1}{\gamma\omega^{r}}(L_{4}+C)+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}.$ (48)
From Lemma 7, Assumption 4, $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, (10), $\textup{dist}(x_{k},X^{*})\leq\delta/2<1/2$ and assuming that $\textup{dist}(x_{k},X^{*})<\varepsilon$, we have
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})||x_{k}-\bar{x}_{k}||^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}||x_{k}-\bar{x}_{k}||+$ (49)
$+||J(x_{k})^{T}F(\bar{x}_{k})||+||J(x_{k+1})^{T}F(\bar{x}_{k})||$
$\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+L_{3}^{r}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|^{1+r}$
@@
Therefore, for $\varepsilon\leq\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r},$ we have $\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*})$. ∎∎
@@
Theorem 3.
Suppose that Assumptions 1-3 and 4 are valid. Let $\{x_{k}\}$ be generated by LMMSS with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$, for all $k$. If $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ as in (50), then $\{dist(x_{k},X^{*})\}$ converges superlinearly to zero. Furthermore, the sequence $\{x_{k}\}$ converges to the solution $\bar{x}\in X^{*}\cap B(x^{*},\delta/2).$
Remark 4.
The main difference from Theorem 2 is that when Assumption 4 holds with $10<r<$ 1 we cannot choose $\lambda_{k}=\|J_{k}^{T}F_{k}\|$ anymore; $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$ is essential for the convergence proof in this case. This becomes clear from the proof of Lemma 6; see inequality (36). In fact, we could choose $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r^{\prime}}$ with $r^{\prime}\leq r$ that all analysis remains valid with minor modifications.
@@ 5 Global Convergence
Now we consider Algorithm 1, a version of LMMSS globalized with line-search for non-zero residue nonlinear least-squares. It is worth to point out that Algorithm 1 differs from the algorithm proposed in [5] in the choice of the LM parameter $\lambda_{k}$ as well as the full step ($\alpha_{k}=1$) acceptance criterion.
Input: $\nu,\zeta,\vartheta,\theta\in(0,1)$, $M>0$ $F$, $J$, $L$, and $x_{0}\in\Omega$.
Set $\lambda_{0}=\left\lVert\nabla\phi({x}_{0})\right\rVert$, $k=0$.
@@
    Compute $d_{k}$, a solution of $(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d=-J_{k}^{T}F_{k}$.
@@
    Choose $m$ as the smallest non-negative integer such that
@@
    Update ${x}_{k+1}={x}_{k}+\alpha_{k}{d}_{k}$, $\lambda_{k+1}=\left\lVert\nabla\phi(x_{k+1})\right\rVert$, $k\leftarrow k+1$,
@@
Algorithm 1 LMMSS with line-search
Initially, we will demonstrate that the sequence of directions $d_{k}$ generated by Algorithm 1 is gradient-related (see [4, Sec. 1.2]). Subsequently, we will establish that any limit point of the sequence produced by this algorithm is a stationary point for (1), regardless of the initial point.
In this section, we assume that Assumption 4 is satisfied with $r=1$ to simplify notation and proofs.
We recall the definition of gradient-related directions from [4, Sec. 1.2].
Definition 1 (Gradient-related).
Let $\{x_{k}\}$ and $\{d_{k}\}$ be sequences in $\mathbb{R}^{n}$. The sequence $\{d_{k}\}$ is said to be gradient-related to $\{x_{k}\}$ if, for each subsequence $\{x_{k}\}_{k\in\mathcal{K}}$ (with $\mathcal{K}\subseteq\mathbb{N}$) converging to a non-stationary point of $\phi$, the corresponding subsequence $\{d_{k}\}_{k\in\mathcal{K}}$ is bounded and satisfies
$\limsup_{k\rightarrow\infty,\ k\in\mathcal{K}}\nabla\phi(x_{k})^{T}d_{k}<0.$ (51)
Proposition 1.
Let $\{d_{k}\}$, $\{x_{k}\}$ be sequences generated by Algorithm 1 starting from $x_{0}\in\mathbb{R}^{n}$. Suppose that Assumption 1 holds in the level set ${\cal L}_{0}=\{x\in\mathbb{R}^{n}\mid\phi(x)\leq\phi(x_{0})\}$. Then, $\{d_{k}\}$ is gradient-related.
Proof.
Suppose that Algorithm 1 generates an infinite sequence $\{x_{k}\}$, i.e., $\|\nabla\phi(x_{k})\|=\|J_{k}^{T}F_{k}\|>0$, for each $k$, then, consequently, $\left\lVert F(x_{k})\right\rVert>0$, and let $\{x_{k}\}_{k\in\mathcal{K}}$, $\mathcal{K}\subseteq\mathbb{N}$, be a subsequence converging to $x_{\infty}$, a non-stationary point for $\phi$, i.e., $\nabla\phi(x_{\infty})\neq 0$. From now on, consider $k\in\mathcal{K}$. From (18) and the definition of $d_{k}$, it follows that
@@
$=-d_{k}^{T}X_{k}^{-T}\left[\begin{array}[]{cc}\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1}d_{k}$
$:=-d_{k}^{T}X_{k}^{-T}\Lambda_{k}X_{k}^{-1}d_{k}.$
By Ostrowski’s inertia law, we have
@@
$\lambda_{n}(X_{k}^{-T}\Lambda_{k}X_{k}^{-1})\geq\gamma\lambda_{n}(\Lambda_{k})\geq\gamma\min\{1,\lambda_{k}\},$
where, in the last inequality, we use the fact that $\sigma_{i,k}^{2}+\mu_{i,k}^{2}=1$. Since $x_{\infty}$ is non-stationary, and $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert$, there exists a $\underline{\lambda}>0$ such that for $k\in{\cal K}$, $\lambda_{k}\geq\underline{\lambda}$. Thus, we obtain
@@
By the continuity of $F(x)$ and $J(x)$ and the convergence of the subsequence $\{x_{k}\}_{\mathcal{K}}$, there exist $L_{J}>0$ and $M_{0}>0$ such that $\|J(x_{k})^{T}F(x_{k})\|\leq M_{0}$ and $\left\lVert J(x_{k})\right\rVert\leq L_{J}$ for each $k$.
Thus, since $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert$, we have
$\|\nabla\phi(x_{k})\|\leq\left(L_{J}^{2}+M_{0}\|L\|^{2}\right)\|d_{k}\|=:M_{1}\|d_{k}\|,\quad\forall k\in{\cal K}.$ (53)
@@
$\limsup_{k\rightarrow\infty,\ k\in\mathcal{K}}\nabla\phi(x_{k})^{T}d_{k}\leq-\frac{\gamma\min\{1,\underline{\lambda}\}}{M_{1}^{2}}\|\nabla\phi(x_{\infty})\|^{2}<0.$
Now we just need to show that $\{d_{k}\}_{k\in\mathcal{K}}$ is bounded. Recall from (19) that
$\|d_{k}\|\leq\|X_{k}\|^{2}\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|,\quad\forall k\geq 0,$ (54)
@@
From Lemma 3, we have $\|X_{k}\|\leq 1/\sqrt{\gamma}$. Now let us analyze the term $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|$. From Lemma 4, it follows that
$\|\Gamma_{k}\|\leq\dfrac{1}{\lambda_{k}},\quad\text{if}\quad 0<\lambda_{k}<1,\quad\text{and}\quad\|\Gamma_{k}\|\leq 1,\quad\text{if}\quad\lambda_{k}\geq 1.$ (55)
@@
    If $10<\lambda_{k}<$ 1, then $\max\{\|\Gamma_{k}\|,1\}\leq\dfrac{1}{\lambda_{k}}$. Since $\lambda_{k}=\|J_{k}^{T}F_{k}\|$, we have
    $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\dfrac{1}{\lambda_{k}}\|J_{k}^{T}F_{k}\|=\dfrac{1}{\|J_{k}^{T}F_{k}\|}\|J_{k}^{T}F_{k}\|=1.$
@@
    If $\lambda_{k}\geq 1$, then $\max\{\|\Gamma_{k}\|,1\}\leq 1$. Thus, $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq M_{0}$.
@@
$\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\max\{1,M_{0}\},\quad\forall k\in{\cal K}.$
@@
$\|d_{k}\|\leq\frac{1}{\gamma}\max\{1,M_{0}\}=:M_{2},\quad\forall k\in{\cal K}.$ (56)
Inequalities (52) and (56) show that for $k\in{\cal K}$, $d_{k}$ is the one determined in Step 2 of Algorithm 1 because it passes the seize of Step 4 for $M=M_{2}$ and $\theta=\gamma\min\{1,\underline{\lambda}\}/M_{1}^{2}$. Thus, the proof is complete. ∎∎
Remark 5.
The above proposition shows that LMMSS directions generated by Algorithm 1 are gradient-related as long as the completeness condition (Assumption 1) holds at every iterate – which also ensures that the solution of (2) is unique. Nevertheless, apart from special cases, e.g when the Jacobian $J(x)$ is full-rank everywhere, we cannot ensure that Assumption 1 holds at every $x_{k}$ generated by Algorithm 1. Hence, we safeguard the algorithm by using the classic LMM direction whenever the LMMSS direction fails to fulfill sufficient conditions for gradient-related property, namely, $\|d_{k}\|\leq M$, for $M>0$ and $-\nabla\phi(x_{k})^{T}d_{k}\geq\theta\|\nabla\phi(x_{k})\|^{p_{1}}$, for $\theta\in(0,1)$ and $p_{1}>0$.
@@
we obtain $-\nabla\phi(x_{k})^{T}d_{k}=\|J_{k}d_{k}\|^{2}+\lambda_{k}\|d_{k}\|^{2}\geq\lambda_{k}\|d_{k}\|^{2},$ and since

### v2

@@ 4.2 Diminishing Rank
In this section, we will examine the scenario where the rank of $J(x_{k})$ decreases as $x_{k}$ approaches the set of stationary points $X^{*}$. In this case, the convergence analysis will depend not only on the upper bound for $\|J(x_{k})^{T}F(\bar{x}_{k})\|$ given in Assumption 3 (or Assumption 4) but also on the specific choices of the LM parameter $\lambda_{k}$.
Lemma 12
Let $x^{*}\in X^{*}$, suppose that Assumptions 1 and 2 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$, and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$. Additionally, assume that Assumption 4 is verified with $r\in]0,1]$, and $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$. If $x_{k},x_{k+1}\in B(x^{*},\delta/2)$ and $\textup{dist}(x_{k},X^{*})<\varepsilon$, where
$\varepsilon\leq\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r}$
with $\eta\in(0,1)$, and $\hat{C}=L_{7}+(1+(1+c_{1})^{1+r})C$, $L_{7}=L_{4}c_{1}^{2}+L_{5}+L_{3}^{r}\left\lVert L\right\rVert^{2}c_{1}$, then
@@
Proof
Note that from Assumption 4 and Lemma 6(a), for $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, we have $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$, with
$c_{1}=\frac{4}{\gamma\omega^{r}}(L_{4}+C)+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}.$ (40)
From Lemma 7, Assumptions 2 and 4, $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, $\textup{dist}(x_{k},X^{*})\leq\delta/2<1/2$ and assuming that $\textup{dist}(x_{k},X^{*})<\varepsilon$, we have
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+L_{3}^{r}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|^{1+r}$
@@
Therefore, for $\varepsilon\leq\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r},$ we have $\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*})$. ∎
@@
Theorem 4
Let $x^{*}\in X^{*}$, suppose that Assumptions 1, 2 and 4 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$ and $r\in]0,1]$, and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$. Let $\{x_{k}\}$ be generated by the LMMSS method with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$ for all $k$, and $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ is given by (41). Then, $\{\textup{dist}(x_{k},X^{*})\}$ converges superlinearly to zero. Moreover, the sequence $\{x_{k}\}$ converges to some $\bar{x}\in X^{*}\cap B(x^{*},\frac{\delta}{2})$.
Remark 4
The main difference from Theorem 3 is that when Assumption 4 holds with $10<r<$ 1 we cannot choose $\lambda_{k}=\|J_{k}^{T}F_{k}\|$ anymore; $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$ is essential for the convergence proof in this case. This becomes clear from the proof of Lemma 6(a); see inequality (31). In fact, we could choose $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r^{\prime}}$ with $r^{\prime}\leq r$ that all analysis remains valid with minor modifications.
To close this section, we analyze local convergence of LMMSS when the Jacobian rank is not constant and Assumption 3 holds.
Lemma 14
Let $x^{*}\in X^{*}$, suppose that Assumptions 1 and 2 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$ and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$. Also, assume that Assumption 3 is verified with
$\sigma<\bar{\sigma}:=\min\left\{(\sigma^{*}_{\min})^{2},\frac{\omega}{(2+(1+\theta\|L\|^{2}))c_{1}}\right\},$
and $\lambda_{k}$, $\theta$ and $\bar{\delta}$ are those in Lemma 6(b). If $x_{k},x_{k+1}\in B(x^{*},\bar{\delta}/2)$, and $\textup{dist}(x_{k},X^{*})<\varepsilon$, where
$\varepsilon\leq\frac{\eta\omega-(2+(1+\theta\|L\|^{2})c_{1})\sigma}{L_{8}},$
with $\eta\in\left]\frac{(2+(1+\theta\|L\|^{2})c_{1})\sigma}{\omega},1\right[$ and $L_{8}=L_{4}(c_{1}^{2}+\theta c_{1}\|L\|^{2})+L_{5}$, then
$\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*}).$
Proof
From Assumption 3 and Lemma 6(b), for $\lambda_{k}$ as in (27), we have $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$, with
$c_{1}=(1/\gamma)\max\{\kappa,(\sigma_{\min}^{*})^{2}/\theta\}+\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}/\sqrt{\gamma}.$ (42)
From Lemma 7, Assumptions 2 and 3, $\textup{dist}(x_{k},X^{*})\leq\delta/2<1/2$ and assuming that $\textup{dist}(x_{k},X^{*})<\varepsilon$, we have
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}||x_{k}-\bar{x}_{k}||$
$+\sigma\|x_{k}-\bar{x}_{k}\|+\sigma\|x_{k+1}-\bar{x}_{k}\|$
$\leq(L_{4}(c_{1}^{2}+\theta c_{1}\|L\|^{2})+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+(2+(1+\theta\|L\|^{2})c_{1})\sigma\|x_{k}-\bar{x}_{k}\|$
$\leq\Big{[}(L_{4}(c_{1}^{2}+\theta c_{1}\|L\|^{2})+L_{5})\varepsilon+(2+(1+\theta\|L\|^{2})c_{1})\sigma\Big{]}\|x_{k}-\bar{x}_{k}\|.$
Now, for $\sigma<\frac{\omega}{(2+(1+\theta\|L\|^{2}))c_{1}}$, for any $\eta\in\left]\frac{(2+(1+\theta\|L\|^{2}))\sigma}{\omega},1\right[$, we have $\eta\omega-(2+(1+\theta\|L\|^{2})c_{1})\sigma>0$. Denoting $L_{8}=L_{4}(c_{1}^{2}+\theta c_{1}\|L\|^{2})+L_{5}$, for $\varepsilon\leq\frac{\eta\omega-(2+(1+\theta\|L\|^{2})c_{1})\sigma}{L_{8}}$, we obtain $\textup{dist}(x_{k+1},X^{*})\leq\eta\textup{dist}(x_{k},X^{*})$. ∎
Again, with the aid of Lemma 14, the proofs of the next results are analogous to those of Lemma 12 and Theorem 4, respectively.
Lemma 15
Suppose that the Assumptions of Lemma 14 are satisfied in $B(x^{*},\delta)$, consider $\bar{\delta}$ from Lemma 6(b) and
$\varepsilon=\min\left\{\frac{\frac{\bar{\delta}}{2}}{1+\frac{c_{1}}{1-\eta}},\frac{\eta\omega-(2+(1+\theta\|L\|^{2})c_{1})\sigma}{L_{8}}\right\}.$ (43)
If $x_{0}\in B(x^{*},\varepsilon)$, then $x_{k+1}\in B(x^{*},\frac{\bar{\delta}}{2})$, and $\textup{dist}(x_{k},X^{*})\leq\varepsilon$ for every $k\in\mathbb{N}$.
Theorem 5
Let $x^{*}\in X^{*}$, suppose that the assumptions of Lemma 14 hold in $B(x^{*},\delta)$. Let $\{x_{k}\}$ be generated by LMMSS with $\alpha_{k}=1$, $\lambda_{k}$ as in (27), for all $k$, $\theta,\bar{\delta}$ from Lemma 6(b). If $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ as in (43), then $\{dist(x_{k},X^{*})\}$ converges linearly to zero. Furthermore, the sequence $\{x_{k}\}$ converges to the solution $\bar{x}\in X^{*}\cap B(x^{*},\bar{\delta}/2)$.
Remark 5
In the results under Assumption 3 that led to Theorem 5, we considered that the Jacobian rank may change in a neighborhood of $X^{*}$. In this case, we emphasize that, in order to apply Lemma 6(b), $\lambda_{k}$ must follow the regime defined by (27), which implies $\sigma\leq\lambda_{k}\leq\theta L_{4}\delta+\theta\sigma$. Thus, in this setting, $\lambda_{k}$ must stays bounded away from zero, as opposed to the previous results in this work.
@@ 5 Global convergence
Now we consider Algorithm 1, a version of LMMSS globalized with line-search for nonzero residual nonlinear least-squares. It is worth to point out that Algorithm 1 differs from the algorithm proposed in Boos2024 in:
  • the choice of the LM parameter $\lambda_{k}$,
  • the full step ($\alpha_{k}=1$) acceptance criterion,
  • a LMM direction as safeguard when LMMSS direction fails to satisfy sufficient descent conditions.
Input: $\nu,\zeta,\vartheta,\xi\in]0,1[$, $M>0$, $F$, $J$, $L$, $\lambda_{0}>0$ and $x_{0}\in\mathbb{R}^{n}$.
Set $k=0$.
@@
    Choose $\lambda_{k}>0$ and compute $d_{k}$, a solution of $(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d=-J_{k}^{T}F_{k}$.
@@
    Choose $m$ as the smallest nonnegative integer such that
@@
    Update ${x}_{k+1}={x}_{k}+\alpha_{k}{d}_{k}$, $k\leftarrow k+1$,
@@
Algorithm 1 Globalized LMMSS
As stated, Algorithm 1 leaves open the choice of the sequence of positive LM parameters $\{\lambda_{k}\}$. Except for the very last theorem of this section, all the results that follow hold for the choices $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r},r\in]0,1]$ or $\lambda_{k}\in[\underline{\lambda},\bar{\lambda}]$, $0<\underline{\lambda}\leq\bar{\lambda}<+\infty$.
[…]
Initially, we will demonstrate that the direction sequence $\{d_{k}\}$ generated by Algorithm 1 is gradient-related to $\{x_{k}\}$ (see (Bertsekas2016, , Eq(1.14))). Subsequently, we will establish that any limit point of the sequence produced by this algorithm is a stationary point for (1), regardless of the initial point.
We recall the definition of gradient-related directions from (Bertsekas2016, , Eq(1.14)).
Definition 2 (Gradient-related)
Let $\{x_{k}\}$ and $\{d_{k}\}$ be sequences in $\mathbb{R}^{n}$ and $\phi:\mathbb{R}^{n}\to\mathbb{R}$ a continuously differentiable function. The sequence $\{d_{k}\}$ is said to be gradient-related to $\{x_{k}\}$ if, for each subsequence $\{x_{k}\}_{k\in\mathcal{K}}$ (with $\mathcal{K}\subseteq\mathbb{N}$) converging to a nonstationary point of $\phi$, the corresponding subsequence $\{d_{k}\}_{k\in\mathcal{K}}$ is bounded and satisfies
$\limsup_{k\rightarrow\infty,\ k\in\mathcal{K}}\nabla\phi(x_{k})^{T}d_{k}<0.$
Remark 6
Let $\{d_{k}\}$, $\{x_{k}\}$ be sequences generated by an iterative algorithm $x_{k+1}=x_{k}+\alpha_{k}d_{k}$, for minimizing $\phi(x)$, where $\alpha_{k}$ fulfills Armijo condition. It is not hard to prove that if $\{d_{k}\}$, $\{x_{k}\}$ satisfy
$\|d_{k}\|\leq\bar{M}\quad\text{and}\quad\nabla\phi(x_{k})^{T}d_{k}\leq-\bar{\xi}\|\nabla\phi(x_{k})\|^{p_{1}},$
for $\bar{M}>0$, $p_{1}>0$ and $\bar{\xi}\in]0,1[$, then $\{d_{k}\}$ is gradient-related.
Proposition 1
Let $\{d_{k}\}$, $\{x_{k}\}$ be sequences generated by Algorithm 1 starting from $x_{0}\in\mathbb{R}^{n}$. Suppose that condition (4) holds in the level set ${\cal L}_{0}=\{x\in\mathbb{R}^{n}\mid\phi(x)\leq\phi(x_{0})\}$, and that either $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$ or $0<\underline{\lambda}\leq\lambda_{k}\leq\bar{\lambda}$. Then, $\{d_{k}\}$ is gradient-related.
Proof
Suppose that Algorithm 1 generates an infinite sequence $\{x_{k}\}$, i.e., $\|\nabla\phi(x_{k})\|=\|J_{k}^{T}F_{k}\|>0$, for each $k$. Then, $\left\lVert F(x_{k})\right\rVert>0$ in this sequence. Let $\{x_{k}\}_{k\in\mathcal{K}}$, $\mathcal{K}\subseteq\mathbb{N}$, be a subsequence converging to $\hat{x}$, a nonstationary point for $\phi$, i.e., $\nabla\phi(\hat{x})\neq 0$. From now on, consider $k\in\mathcal{K}$. From (13) and the definition of $d_{k}$, it follows that
@@
$=-d_{k}^{T}X_{k}^{-T}\left[\begin{array}[]{cc}\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1}d_{k}$ (45)
$:=-d_{k}^{T}X_{k}^{-T}\Lambda_{k}X_{k}^{-1}d_{k}.$ (46)
By Ostrowski’s inertia law (ostrowski, , Theorem 1), we have
@@
$\lambda_{n}(X_{k}^{-T}\Lambda_{k}X_{k}^{-1})\geq\gamma\lambda_{n}(\Lambda_{k})\geq\gamma\min\{1,\lambda_{k}\},$ (47)
where, in the last inequality, we use the fact that $\sigma_{i,k}^{2}+\mu_{i,k}^{2}=1$.
If $\lambda_{k}$ is chosen such that $\lambda_{k}\geq\underline{\lambda}>0$, from (46) and (47) we obtain
@@
If $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, (48) also holds, because $\left\lVert J_{k}^{T}F_{k}\right\rVert>0$, for $k\in{\cal K}$, and $\hat{x}$ is nonstationary, which implies the existence of $\underline{\lambda}>0$ such that for $k\in{\cal K}$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|\geq\underline{\lambda}$.
@@
By the continuity of $F(x)$ and $J(x)$ and the convergence of the subsequence $\{x_{k}\}_{\mathcal{K}}$, there exist $L_{J}>0$ and $M_{0}>0$ such that $\|J(x_{k})^{T}F(x_{k})\|\leq M_{0}$ and $\left\lVert J(x_{k})\right\rVert\leq L_{J}$ for each $k\in{\cal K}$.
Thus, if either $0<\underline{\lambda}\leq\lambda_{k}\leq\bar{\lambda}$ or $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}\leq M_{0}^{r}=:\bar{\lambda}$, we have
$\|\nabla\phi(x_{k})\|\leq\left(L_{J}^{2}+\bar{\lambda}\|L\|^{2}\right)\|d_{k}\|=:M_{1}\|d_{k}\|,\quad\forall k\in{\cal K}.$ (49)
@@
$\limsup_{k\rightarrow\infty,\ k\in\mathcal{K}}\nabla\phi(x_{k})^{T}d_{k}\leq-\frac{\gamma\min\{1,\underline{\lambda}\}}{M_{1}^{2}}\|\nabla\phi(\hat{x})\|^{2}<0.$
Now we just need to show that $\{d_{k}\}_{k\in\mathcal{K}}$ is bounded.
Recall from (14) that
$\|d_{k}\|\leq\|X_{k}\|^{2}\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|,\quad\forall k\geq 0,$
@@
From Lemma 3, we have $\|X_{k}\|\leq 1/\sqrt{\gamma}$. Now let us analyze the term $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|$.
From Lemma 4, it follows that
$\|\Gamma_{k}\|\leq\dfrac{1}{\lambda_{k}},\quad\text{if}\quad 0<\lambda_{k}<1,\quad\text{and}\quad\|\Gamma_{k}\|\leq 1,\quad\text{if}\quad\lambda_{k}\geq 1.$
@@
    If $10<\lambda_{k}<$ 1, then
    $\max\{\|\Gamma_{k}\|,1\}\leq\dfrac{1}{\lambda_{k}}.$
    If $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$, we have
    $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\dfrac{1}{\lambda_{k}}\|J_{k}^{T}F_{k}\|=\dfrac{1}{\|J_{k}^{T}F_{k}\|^{r}}\|J_{k}^{T}F_{k}\|=\|J_{k}^{T}F_{k}\|^{1-r}.$
    If $\lambda_{k}\geq\underline{\lambda}>0$, we obtain
    $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\dfrac{1}{\lambda_{k}}\|J_{k}^{T}F_{k}\|=\dfrac{1}{\underline{\lambda}}\|J_{k}^{T}F_{k}\|.$
    In both cases the right hand side is bounded by $\max\left\{M_{0}^{1-r},M_{0}/\underline{\lambda}\right\}$.
@@
    If $\lambda_{k}\geq 1$, then $\max\{\|\Gamma_{k}\|,1\}\leq 1$. Thus,
    $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\|J_{k}^{T}F_{k}\|\leq M_{0},\quad\forall k\in{\cal K}.$
@@
$\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\max\{M_{0}^{1-r},M_{0}/\underline{\lambda},M_{0}\}=:\bar{M}_{0},\quad\forall k\in{\cal K}.$
@@
$\|d_{k}\|\leq\frac{1}{\gamma}\max\{\bar{M}_{0},M_{0}\}=:M_{2},\quad\forall k\in{\cal K}.$
Thus, the proof is complete. ∎
The above proposition shows that directions $d_{k}$ generated at Step 2 of Algorithm 1 are gradient-related, for both choices $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$, or $0<\underline{\lambda}\leq\lambda_{k}\leq\bar{\lambda}$, as long as the completeness condition (4) holds at every iterate. Nevertheless, apart from special cases, e.g when the Jacobian $J(x)$ is full-rank everywhere, we cannot ensure that such condition holds at every $x_{k}$ generated by Algorithm 1 starting at an arbitrary $x_{0}\in\mathbb{R}^{n}$. Hence, we safeguard the algorithm by using the classic LMM direction whenever the LMMSS direction fails to fulfill sufficient conditions for gradient-related property, namely, those of Remark 6.
@@
we obtain
$-\nabla\phi(x_{k})^{T}d_{k}=\|J_{k}d_{k}\|^{2}+\lambda_{k}\|d_{k}\|^{2}\geq\lambda_{k}\|d_{k}\|^{2},$
and since